### v1

# Designing for Distributed Heterogeneous Modularity: On Software Architecture and Deployment of the MoonBots

Abstract
This paper presents the software architecture and deployment strategy behind the MoonBot platform: a modular space robotic system composed of heterogeneous components distributed across multiple computers, networks and ultimately celestial bodies. We introduce a principled approach to distributed, heterogeneous modularity, extending modular robotics beyond physical reconfiguration to software, communication and orchestration. We detail the architecture of our system that integrates component-based design, a data-oriented communication model using ROS2 and Zenoh, and a deployment orchestrator capable of managing complex multi-module assemblies. These abstractions enable dynamic reconfiguration, decentralized control, and seamless collaboration between numerous operators and modules. At the heart of this system lies our open-source Motion Stack software, validated by months of field deployment with self-assembling robots, inter-robot cooperation, and remote operation. Our architecture tackles the significant hurdles of modular robotics by significantly reducing integration and maintenance overhead, while remaining scalable and robust. Although developed with space in mind, we propose generalizable patterns for designing robotic systems that must scale across time, hardware, teams and operational environments.
I Introduction
Achieving true modularity in robotics would mark a significant leap forward [1, 2], especially in the context of space exploration [3, 4, 5]. Most robotic systems today are engineered as monolithic, task-specific platforms, requiring the task or environment to adapt to the robot rather than the opposite [6, 7, 8]. This rigid coupling between form and function significantly limits adaptability.
[…]
Figure 1: A variety of MoonBot assemblies. (a): Dragon assembly – 4 modules, made of 2 MoonBot G-line limbs and 2 MoonBot H-line V1 wheels. Specialized in nimble navigation. (b): Dragon x4 assembly – 8 modules, made of a mix of MoonBot H-line V1 and V2, wheels and limbs. Specialized in redundancy. (c): Tricycle assembly – 6 MoonBot H-line V1 modules. Spacialized in cargo. (d): MoonBot H-line V1 limb and V2 wheel assembling themselves.
As a concrete example, no single robot can palletize boxes on an assembly line, drive on public roads, and traverse mountainous terrain. While excellent technologies exist for each task individually [9, 10, 6], they remain siloed and cannot be unified into a single, adaptive system. Humans, in contrast, excel at switching between domains – a capability that continues to inspire the pursuit of general-purpose humanoid robots [11]. But, as illustrated in Fig. 1, modular robots offer an alternative vision: breaking the monolith to build reconfigurable systems capable of adapting to the task, environment, and mission [4, 1, 2].
Our project: MoonShot [12], envisions a fleet of MoonBot robots building lunar infrastructures. As illustrated Figs. 1 and 2 and described in our article [13], MoonBots reconfigure in a multitude of shapes, specialized in diverse tasks – exploration, mapping, construction, and repair – and distributed over wide networks. Modularity is key for such space projects [5, 14]: Redundancy is improved with replaceable parts; weight is reduced by sending modules instead of specialized robots; load is balanced between servers; the system can be upgraded through the years [3, 4].
To proceed, we must clarify our broader notion of modularity. While existing robotic frameworks define it through physical reconfiguration [15, 16], our approach distinguishes itself by not tying modularity to physical shape. In the literature we find three emergent family of modularity: “graft modularity” (augmenting capability by adding components [7, 8]), “clone modularity” (scaling via identical units [2, 17]), and “swarm modularity” (emergent behavior from many similar agents [10, 18]). Our framework aim for a superset, novel to robotics, encompassing all those approaches: “distributed heterogeneous modularity,” where modules differ in form, function, and communication, and are distributed across both physical and computational infrastructure.
We can find our large-scale modularity practiced in space systems [3]. The International Space Station exemplifies this mindset with heterogeneous physical modules, but also distributed components across orbit and ground, cooperating for control, analysis, and fault recovery.
Distributed heterogeneous modularity is now widespread in modern software, with Internet of Tings (IoT) and cloud systems [19]. These domains have developed mature, scalable patterns for managing complexity across heterogeneous layers of infrastructure. We observe those technologies – DDS, MQTT, and Zenoh [20, 21, 22] – arriving in the field of robotics, giving birth to Industrial IoT (IIoT) [19]. Based on those ideas ROS2 [23] is now largely unavoidable in robotics, successfully implementing distributed computation, standardized messaging and packaging. We will see that ROS2 provides most – but not all – of the solutions we need.
Figure 2: Zenoh network overview of distributed Moon bases and robot assemblies. It combines multiple topologies to maximize reliability and performance under various communication interfaces.
This paper describes the software architecture required to support the MoonShot project: enabling modularity across form, function, and location. Our system must scale in both hardware and network dimensions, supporting reconfigurable robot assemblies as well as the interplanetary communication patterns envisioned in Fig. 2. Furthermore, this being at the laboratory experiment stage, other requirements take precedence over space consideration. Notably development, prototyping, developer onboarding should be simple, and the architecture should be robust to frequent changes. A keen reader will notice that this is also analogous to modularity.
II Method
II-A Design Philosophy
As highlighted by Conway [24], product and software architecture converges towards the structure of the group of people making it. In robotics, we find that the software is highly correlated to the shape of the robot.
For us, this creates a critical contradiction, because a deeply modular robot should not assume a shape. Hence, the developer must actively fight against this pull toward a fix architecture and inappropriate assumptions.
To avoid such shortcomings, we propose a strong component-based software approach [25]. We will avoid encoding relationships between software components, because this would imply relationship between people and robot shape. Instead, we treat the system as assemblies of loosely-coupled components, where behavior emerges from data.
Core to this model is separation of concerns [26, 25]. Components must not be aware of the internal logic – or even existence – of other components. They should operate solely on resources, exposed via a medium of data. Our architecture should be shaped by data, not by the shape of the robot. It is then, the responsibility of higher-order components – solvers, planners, orchestrators, and AI agents – to compose behavior by acting on those data resources.
II-B Network
As in industrial IoT (IIoT) systems, the network is a critical enabler of modularity [19, 22, 17]. Given our goal Fig. 2 and philosophy Sec. II-A, we require a data transport layer offering strong separation of concerns, while supporting long range and unreliable communications.
Traditional Data Delivery Systems (DDS) are efficient and relatively simple to deploy [20]. However, they are strictly peer-to-peer, thus violate separation of concerns by tightly coupling data flow between all participants.
In contrast, MQTT communicates through a central broker [21]. While this decouples participants, it introduces a single point of latency, failure, and bandwidth constraint.
Zenoh [22] offers the best of both: peer-to-peer, brokered, and mesh communication can coexist. Hence, the network topology can adapt to the mission and robot. Independent software routers manage routing and optimization, making network orchestration the sole responsibility of Zenoh.
Fig. 2 illustrates powerful concepts. Wireless communication links are routed through Zenoh routers, isolating fragile and long-range connections. Meanwhile, high-performance peer-to-peer links are used on intra-base and intra-robot communication. Redundancy naturally emerges through multiple radio paths and fallback wired links within each robot assembly (e.g., between limb #1, #2 and wheel #1, #2).
Due to hardware constraints, wired communication between modules is not yet available. All inter-module traffic is currently routed over Wi-Fi. Although a brittle, worst-case scenario, this setup is common in many robotic systems.
II-C Launch and Orchestration
As the number of robot modules grows – each potentially composed of multiple onboard computers – the task of launching the correct processes, with appropriate configuration, on the correct machines becomes exponentially more complex, time-consuming, and error-prone. Our orchestrator [25] addresses this by reducing operator intervention down to three actions per robot: update, build, and start. While not fully automated, this already provides critical reliability and speed during deployment.
Additionally, orchestration encodes static information that is not transmitted over the Zenoh network. Encoding every behavior as runtime messages would lead to unnecessary complexity and fragility.
Our orchestrator consists of the following steps:
  1. 1.
    Version Control: Updates numerous private repositories to the correct branches for the current experiment.
  2. 2.
    Updater and Build: Resolves dependencies, then uses the Doit build system [27] to resolve a graph of offline installation and compilation tasks.
  3. 3.
    Robot Assembler: A Python system that defines the software components of every module for use in a complete assembly. It deduces assembly-specific configurations (e.g., kinematic chains, URDF), computer-specific settings (e.g., executable paths, joints under control), and hardware-specific parameters (e.g., calibration data, motor type).
  4. 4.
    Component Launcher: Identifies the host machine, launching the appropriate components. Execution is remotely started and internally delegated to the ROS2 launch system.
II-D Components
Our component structure, shown in Fig. 3 and detailed in Tab. I, is designed to minimize robot-specific code and enforce modularity down to the subcomponent level.
TABLE I: Roles and scopes of subcomponents Fig. 3.
Block Role Specific to Notes
Core Main logic None Pure python
Injection Extend the core Core type Tested, tunable
Overrides Alter the core Robot System-specific
API Simple interface None Beginner friendly
Interface Bridge Core type ROS2 wrapper
Executor Call the core OS ROS2 executor
Comm. Data Transport Architecture ROS2 over Zenoh
Figure 3: Structure of a software component, pushing for separation of concern down to the subcomponent level.
Modular robotics present numerous hardware and system-specific variations. To accommodate each specificity, the code often branches with multiple execution paths, ultimately becoming difficult to maintain.
Our model avoids this by isolating robot-specific logic within the well-defined Injection and Override blocks [26, 25] – both of which are chosen by the orchestrator (Sec. II-C). For example, Injection may be a telemetry system, while Override handles a broken joint.
We deliberately avoid using launch parameters for hardware-specific options. First, Core developers cannot anticipate all possible customization points. Second, launch parameters are limiting and static, thus unsuitable for runtime behavior. Third, following separation of concerns, the launcher should only expect to interact with the original Core class. While tightly integrated, the orchestrator aims to be an independent system.
The highly reusable API subcomponent abstracts low-level communication and safety-critical logic behind a consistent interface. As shown in our previous work [28], actuator synchronization can be robustly implemented without requiring detailed system knowledge – allowing safe operations on any actuators.
The strength of this whole structure lies in its composability: each subcomponent can be reused, swapped, or tuned with minimal branching and coupling to other systems.
II-E Implementation and Test Campaigns
The system was implemented on the MoonBots (Fig. 1) and PCs of collaborators interacting with the robots. ROS2 serves as our backbone, providing both DDS and Zenoh compatibility. The MoonBots and user PCs run Ubuntu 22.04 or 24.04 depending on hardware compatibility. Although ROS2 does not fully meet our separation of concerns requirement – due to its tight coupling between operating system of all participants – the need for rapid prototyping and onboarding motivated its adoption.
The system was tested and improved gradually by intensive test campaigns: three field tests totaling nine weeks at the JAXA sand field [29]; a five-day experiment in DLR’s LUNA lunar simulation facility (Fig. 1a) [30]; and a six-day campaign at the Osaka 2025 Universal Exposition, with daily operations of ten hours. Referring to Fig. 2, these tests implemented MoonBase #1, where multiple students simultaneously operated multiple robots locally. Some tests also involved remote operations, simulating a lunar base controlling robots from afar. In one proof-of-concept trial at the Osaka Exposition, we demonstrated precise teleoperation across multiple wireless links: Operator $\rightarrow$ WiFi $\rightarrow$ Internet $\rightarrow$ Cellular 5G $\rightarrow$ WiFi $\rightarrow$ Robots.
Our primary goal in those campaign was to provide a flexible platform for all our collaborators to control the robots. Those experiments provided us with critical feedback, observation and data (rosbag), shaping the current system architecture.
III Results
III-A Modularity
Our system achieves broad modularity. We curretnly support seven robot modules across two hardware families (H-line and G-line). Importantly, those family lines have major hardware differences – number of degree-of-freedom, motor controller, embedded computer – reflected in software by changes to the Injection and Override blocks Fig. 3. This diversity shown Figs. 1 and 4 is comprehensively managed by the orchestrator, allowing for mixed configurations.
[…]
Figure 4: Two assemblies of independent robot modules, connecting to form one large assembly. (a) Left: Moonbot Dragon assembly, made of 2 H-line wheels V2 and 2 H-line limbs V2. Right: Variant with V1 modules. (b) The large heterogeneous assembly moving.
As outlined in Sec. I, modularity must extend beyond hardware to the full robotic system. During experiments involving two Dragon assemblies (Fig. 4.a) multiple ground control teams operated the robots concurrently from separate machines. Tab. II details the 60 components simultaneously involved. Each robot and control module acted independently, communicating solely via data – enabling decentralized operation with minimal coupling.
Operator collaboration emerged naturally, with seamless task sharing across independent component. In Fig. 4, coordinated steering was achieved as two human operators controlled the front and rear halves – demonstrating the system’s decentralized and composable behavior in practice.
TABLE II: Components during the distributed operation of the Dragon assemblies in Fig. 4. (GC: Ground Control)
Component Count Location
Motor interface 8 Robot
Joint Manager 8 Robot
Kinematics Manager 4 Robot
Health monitor 8 Robot
Location publisher 16 Robot
Human operator 3 GC
Autonomous operator 6 GC
Data monitoring 5 GC
Data archiving 2 GC
III-B Development Time
TABLE III: Estimated time required for common development and deployment tasks under our system.
Action Time estimate
Adding a new robot family line 1-2 week
Adding a new module to a family 2 days
Adding an updated module revision 1 day
Coding a new assembly template 1 day
Swapping robots of an assembly 20 seconds
Getting an assembly’s parameters & URDF Instantaneous
Entire software instalation 30 minute
Switching between sim and real Instantaneous
Starting a remote robot 1 minute
Learning ROS2 2 weeks
Onboarding a new developer 1 week
Learning only the Python API 2 day
Starting one Zenoh router 1 minute
Applying communication layer changes 1 minute
The component-based architecture substantially reduced development effort by focussing on modularity. This proved critical during field-testing where experiments are conducted by many students and robots often replaced. The estimated durations of the common operations are outlined Tab. III.
System Integration. Adding or modifying robot families, modules, or assemblies requires minimal changes – mostly confined to the Injection, Override, and metadata layers. The orchestrator automates system configuration and URDF generation, enabling rapid prototyping and reuse.
Deployment and Simulation. Setting up a full robot environment is fast and reproducible with the orchestrator. Simulation is deeply integrated: hardware components are automatically replaced with virtual counterparts when orchestration detects a non-robot host. This makes sim-to-real transitions instantaneous and consistent.
Onboarding and Productivity. The Python API dramatically lowers the barrier to entry. Developers can interact with components without needing to understand low-level ROS2 concepts, which reduces onboarding time and did prevent many hardware damages.
Communication Reconfiguration. The separation of communication logic enables fast, non-intrusive changes to network transport. These changes can be deployed quickly – critical during large experiments involving many end-points.
III-C ROS2 Limitations
While ROS2 is a backbone of our system, its node-centric design conflicts with our data-oriented model. The ROS2 ecosystem encourages minimizing the number of nodes and data endpoints (e.g., publishers and subscribers). Additionally, all endpoints contribute to a single, globally shared “rosgraph”. This tightly couples components and breaks separation of concerns.
This is especially evident in the handling of transforms: all transforms in ROS2 are traditionally aggregated on a global /tf topic. As a result, an Earth station that only needs the pose of a single robot must listen to the entire – interplanetary – transform stream. Data aggregation, batching, routing and stream optimization should instead be delegated to the network, by keeping data simple and decoupled.
Beyond architectural concerns, this introduces excessive network traffic, processing, and fragility. These limitations are not only theoretical, as will demonstrate the next section.
Finally, the incompatibilities between ROS2 versions – tied to very specific and heavy distributions of Ubuntu – is a major drawback for large scale collaboration. Teams, components and robots become incompatible and fragmented because of requirements unrelated to the project.
III-D Network
TABLE IV: Comparison of network behavior between DDS and Zenoh, measured on one robot with 6 other active robots.
DDS Zenoh
Component startup 10 s to 1 min 1 s
Startup bandwidth 69 Mb/s 1.65 Mb/s
Startup stutter 20 s 1s
Average latency 1.6 ms 2.2 ms
Average transmitted bandwidth 1,200 Kb/s 1,130 Kb/s
Average received bandwidth 550 Kb/s 250 Kb/s
Peak bandwidth 69 Mb/s 2 Mb/s
Connectivity recovery time 20s 2s
Max stable robot count 4 10+
We began development using ROS2’s default DDS backend, improved by discovery servers. However, DDS quickly suffered from severe startup and recovery overheads. The analysis of field test’s data (rosbags) highlights those problems Tab. IV. Most problematic being that not only ROS2, but network-wide stutters happen using DDS.
These issues persisted despite using a high-end Wi-Fi 7 setup (multi-link operation, dedicated APs, isolated traffic) – conditions more favorable than most real-world deployments. Still, DDS was not reliable enough to support more than four robots transmitting only basic joint and state information.
As anticipated in Sec. II-B, DDS’s peer-to-peer design fails to scale under our distributed, modular architecture.
Switching to Zenoh as ROS2 communication middleware resolved these issues. The clean separation in our architecture (Secs. II-C and 3) allowed us to migrate within a day. With Zenoh, reactiveness improved, bandwidth stabilized, and collaboration across many components (Tab. II) became seamless – allowing for reliable, large assemblies (Fig. 4). Zenoh ultimately allowed seamless remote operations, notably the Osaka trial previously explained Sec. II-E.
The startup stutter is however still slightly present with Zenoh, but it is short and only affects ROS2 communications.
IV Discussion
IV-A Future Transport Layer Direction
While Zenoh is already our ROS2 communication middleware, its full benefits are constrained by ROS2– as preempted by Sec. II-E. A full transition to pure Zenoh would resolve many architecture problems highlighted Sec. III-C:
  • Only data flows, reducing overhead and brittleness.
  • Lightweight deployment: Low system requirements.
  • Scalability: Thousands of data sources and sinks.
  • Advanced routing: data is precisely managed and monitored by the routers.
  • Storage: Native support for distributed data storage (e.g., telemetry, LiDAR scans, video logs).
We are actively exploring a full Zenoh implementation to better match our data-oriented, distributed model. However, this comes with ecosystem tradeoffs: ROS2 tools would need reimplementation, collaborators would need to learn new skills, and compatibility with the broader ROS2 community would be lost. As such, our near-term goal is a pure Zenoh prototype, then a compatibility bridge to ROS2.
Finally, future work will include testing under degraded conditions resembling space networks – high-latency, low-bandwidth, and intermittent links.
IV-B Toward Smarter Orchestration
Currently, the orchestrator relies on a manual launch procedure: an operator specifies the assembly and initiates the launch. While manual, it is infrequent and human safety assessment is essential – especially in high-stakes space robotics. Furthermore, the assembly information takes less than 20 seconds to specify via the robot-assembler detailed Secs. II-C and III.
We believe that full automation of the orchestrator offers limited practical value. Automatically detecting assemblies would require each robot to identify itself, introducing coupling and added complexity.
A more beneficial improvement would be enabling local orchestrators to synchronize with a central database. Currently, they operate fully offline and are manually updated on each robot – which is error-prone and time-consuming. Network-based synchronization would improve maintainability without compromising modularity or safety.
IV-C Applicability
This system has not yet been tested under space-like constraints. Our hardware uses standard consumer components, and our communication relies on terrestrial cellular network, Wi-Fi and LAN, rather than space-qualified radios. This is by design: the primary objective of this work is not a space-specific solution, but to discover architectural guidelines applicable to future modular and distributed robotic systems.
Our goal is to surface principles that scale across all domains through our prototypes. The lessons presented throughout the paper – component separation (Secs. II-A, II-D and III-A), data-centric communication (Secs. II-B and III-D), and orchestrated deployment (Sec. II-C) – hold beyond space applications.
Furthermore, our most fundamental components – joint manager, kinematic manager, API [28], operator interface – are available open-source in our Motion Stack software (https://github.com/2lian/Motion-Stack).
Ultimately, we advocate for software structures that are resilient to change: of robot shape, team composition, communication infrastructure, or mission scope.
V Conclusion
This work presents a software architecture and deployment strategy for distributed, heterogeneous modularity in robotics. Building on principles of separation of concerns, component-based design, and data-centric communication, we demonstrated a system capable of operating large, reconfigurable assemblies composed of independently designed and controlled modules.
The modularity, flexibility and network functionality of our architecture was improved and validated through extensive field-testing, while maintaining low development overhead. Key enablers include the orchestrator (Sec. II-C), which automates deployment across diverse modules; the clean component structure (Sec. II-D); and the Zenoh-based communication backbone (Sec. II-B), which resolves limitations inherent in traditional ROS2 and DDS systems (Secs. III-C and III-D).
While not yet space-hardened, our system was deliberately with generality in mind (Sec. IV-C). Our architecture aims to be agnostic to physical implementations, and thus applicable to industrial, terrestrial, and interplanetary deployments alike. Future work will address space-specific constraints, further promising Zenoh tooling, and improving orchestrator via centralized server. This contribution serves as a practical reference for designing scalable, modular robotic systems.
More broadly, we advocate for adaptive, distributed deployments as a foundation not just for modular space robotics, but for the future of robotics as a whole.